OpenStudy (anonymous):

\[ax^2 + b \left[ x \right] + c =0\] the sum of roots is?

OpenStudy (shubhamsrg):

why is that x in brackets? is it supposed to be mod ?

OpenStudy (anonymous):

yes it is mod

OpenStudy (apoorvk):

matlab, for x>0 \(ax^2 + bx + c =0\) and, for x<0 \(ax^2 +bx + c=0\) so for the first one, sum of roots = -b/a and for the second one, b/a

OpenStudy (anonymous):

thxx so the answer is 0

OpenStudy (shubhamsrg):

0 ?? nops..it will have 2 ans for 2 diff cases..

OpenStudy (apoorvk):

Nope, they are just two different cases defining equation.